OpenStudy (anonymous):

Help will give medal and fan whoevr answer my q.... Samantha uses her credit card to pay for a collision repair that ended up totaling $1,433.41. She can pay off up to $500 a month. The card has an annual rate of 16.5%. How much will she pay in interest? $236.51 $472.64 $47.24 $39.23

OpenStudy (anonymous):

$39.23

OpenStudy (anonymous):

@fernando22 how did u get tht

OpenStudy (anonymous):

TIP:/ Make sure you divide the the annual rate by twelve before applying compound intrest formula. M = P( 1 + i )n M is the final amount including the principal. P is the principal amount. i is the rate of interest per year. n is the number of years invested.

OpenStudy (anonymous):

@Henry.Lister .. u confused me

OpenStudy (anonymous):

So in this case P=1433.41, i=16.5, n=(1/4).

OpenStudy (anonymous):

Or I = 0.165 not 100% sure on that.

OpenStudy (anonymous):

There is something called a compound intrest formula which is what you pay on a credit card. You plug in the data, and voila you have a result.

OpenStudy (anonymous):

so its 39.23?

jimthompson5910 (jim_thompson5910):

Month 1: I = PRT I = 1433.41*0.165*(1/12) I = 19.71 So at the end of the first month, she's paying $19.71 in interest. The rest of the payment is principal (so the principal of the first monthly payment is 500 - 19.71 = 480.29) So the remaining balance is 1433.41 - 480.29 = 953.12 --------------------------------------------------------------------- Month 1: I = PRT I = 953.12*0.165*(1/12) I = 13.11 So at the end of the second month, she's paying $13.11 in interest. The rest of the payment is principal (so the principal of the second monthly payment is 500 - 13.11 = 486.89) So the remaining balance is 953.12 - 486.89 = 466.23 ------------------------------------------------------------------- Month 3: The remaining balance is $466.23, which is less than $500, so she pays no interest this month. So to sum things up, she paid $19.71 in interest for month 1 and $13.11 for month 2 In total, she paid 19.71 + 13.11 = 32.81 dollars in interest I may have made an error, but this makes sense to me at least, but I could be off.

OpenStudy (anonymous):

anyone ther?

OpenStudy (anonymous):

@jim_thompson5910 thank u for ur effort

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

nayone else want to take a shotat it

jimthompson5910 (jim_thompson5910):

np, I'm missing around $6 in interest which is puzzling because I have a calculator which is giving me 32.81 dollars in interest

OpenStudy (anonymous):

$236.51